What are the terms ao, a, a2, and a3 of the sequence (an, where an equals a) 2+1? c) In/2]? b) (n1)+1? d) In/2] In/21?
Answer #1

a) a(0)=2^0+1=2, a(1)=2^1+1=3 a(2)=2^2+1=5 a(3)=2^3+1=9

b)a(0)=(0+1)^(0+1)=1, a(1)=(1+1)^(1+1)=2^2=4 a(2)= (2+1)^(2+1)=3^3=27 a(3)=(3+1)^(3+1)=4^4=256

c)[n/2] where[ ] is the lower interger which is less than n/2, thus  a(0)=[0/2]=0, a(1)=[1/2]=[0.5]=0

a(2)= [2/2]=[1]=1 a(3)=[3/2]=[1.5]=1

d) Here we have lower integer as well as higher integer(floor and ceiling), thus

a(0)=[0/2]+[0/2]=0 a(1)=[1/2]+[1/2]=0+1=1 a(2)=[2/2]+[2/2]=1+1=2 a(3)=[3/2]+[3/2]=[1.5]+[1.5]=1+2=3